K. Baggett, J. Benesch, M. Bickley, T. Carlino, G. Ciovati, J. Creel, E. Daly, K. Davis, M. Drury, C. Hovater, N. Huque, F. Humphrey, T. Ganey, C. Ginsburg, A. Kimber, L. King, G. Krafft, F. Marhauser, R. Michaud, C. Mounts, R. Nelson, W. Oren, A. Palczewski., T. Plawski, T. Powers, J. Preble, C. Reece, T. Reilly, A. Seryi, A. Solopova, M. Spata, and P. Vasilauskis
Agenda and Notes
1. Operations Status and Summary/Review (Bickley/Powers)
Resume every other week meeting on RF/SRF issues.
Quenches observed in 2L26-8 over weekend. Hard quenches indicated from analysis of field data.
Not able to recover cavity, even at low gradient without high heat load?
Correlation with end of Linac and Differential pumping posited and discussed.
2. Plasma Processing Update (Powers)
3. Progress on Particle Counter (Geng/Spata)
Particle Counter getting closer. Now design is focusing on "laser blanket" for aperture design. Detect phase shifts in interferometer arrangement.
July 22 updated solution and updated cost presented to division leadership. Waiting for favorable financial decision.
New Ion pump evaluation;